# Graph-theoretical estimates of the diameters of the Rubik’s Cube groups

### v1

@@ IV.5 The 3$\times$3$\times$3 Cube in the quarter-turn metric
Figure 6: A partial Cayley graph of the 3$\times$3$\times$3 Rubik’s Cube group in the quarter-turn metric, which is symmetric with degree ($k$) 12. The order ($n$) is $4.33\times 10^{19}$, the girth ($g$) is 4, the diameter ($d$) is 26, and the number of cycles with length $g$ per vertex ($\eta$) is 18.
Figure 6 is the local (up to distance 2) of the Cayley graph of the 3$\times$3$\times$3 Rubik’s Cube group in the quarter-turn metric.Hirata (2024) It is a symmetric graph with $k=12$, $g=4$, and $\eta=18$. The order of the graph is $n=4.33\times 10^{19}$ according to combinatorial logic.Hofstadter (1981); Singmaster (1981); Joyner (2008); Hirata (2024) The order of this group/graph is so high that an explicit enumeration of distance array $\bm{d}$ is nearly impossible. Nonetheless, by their Herculean computations, Rokicki et al.  determined the diameter to be 26.Rokicki et al. (2013, 2014a, 2014b)
Equation (LABEL:dmin1) places the lower bound for $d$ at $d_{\text{min}}=20$, whereas the probabilistic estimation [Eq. (21)] suggests $d_{\text{probab}}=24.8$. They are, respectively, consistent with and in reasonable agreement with the actual diameter of 26.
@@ V Conclusions
Precisely determining the diameter of a large graph of the order $n$ in excess of millions or billions poses a major computational challenge. When $n$ can be known by a combinatorial or other argument and furthermore the graph has a high symmetry with its local parameters easily determined by inspection, one may estimate the diameter using an algebraic, geometrical, or probabilistic argument. In our previous study, we proposed a probabilistic estimate of the diameter for a large Cayley graph, which reproduces the diameters of the Rubik’s Cube groups within 2 of the actual values or often exactly. Here, we introduced a strict lower bound for the diameter for a symmetric Cayley graph, calculable with $n$ and other easily determined local parameters such as the degree $k$, girth $g$, and number $\eta$ of cycles with length $g$ per vertex. Similar, but more crude estimates of the diameters of the Rubik’s Cube groups have been made,Singmaster (1981) but this study is distinguished from them for presenting a rigorous lower bound also applicable to other symmetric graphs. When applied to symmetric Cayley graphs of some Rubik’s Cube groups, it yields reasonably tight lower bounds for the diameters, which are 60% to 77% of the correct values for large-$n$ graphs.

### v2

@@ IV.5 The 3$\times$3$\times$3 Cube in the quarter-turn metric
Figure 6: A partial Cayley graph of the 3$\times$3$\times$3 Rubik’s Cube group in the quarter-turn metric, which is symmetric with degree ($k$) 12. The order ($n$) is $4.33\times 10^{19}$, the girth ($g$) is 4, the diameter ($d$) is 26, and the number ($\eta$) of $g$-cycles per vertex is 18.
Figure 6 is the local (within distance 2) Cayley graph of the 3$\times$3$\times$3 Rubik’s Cube group in the quarter-turn metric.Hirata (2024) It is a symmetric graph with $k=12$, $g=4$, and $\eta=18$. The order of the graph is $n=4.33\times 10^{19}$ according to combinatorial logic.Hofstadter (1981); Singmaster (1981); Joyner (2008); Hirata (2024) The order of this group/graph is so high that an explicit enumeration of distance array $\bm{d}$ is nearly impossible. Nonetheless, by their Herculean computations, Rokicki et al.  determined the diameter to be 26.Rokicki et al. (2013, 2014a, 2014b)
Equation (LABEL:dmin1) places the lower bound for $d$ at $d_{\text{min}}=20$, whereas the probabilistic estimation [Eq. (21)] suggests $d_{\text{probab}}=24.8$. They are, respectively, consistent with and in reasonable agreement with the actual diameter of 26. The $d_{\text{min}}=20$ of this work is slightly tighter lower bound than the $d_{\text{min}}^{\text{BV}}=19$, which is understandable because the former uses more information about the graph such as its symmetry and girth. On the other hand, $d_{\text{max}}^{\text{BV}}=23$ undershoots the diameter, indicating that $n=4.33\times 10^{19}$ is still far from the asymptotic limit and the use of $\epsilon=0$ is inappropriate.
@@ V Conclusions
Precisely determining the diameter of a large graph of the order $n$ in excess of millions or billions poses a major computational challenge. When $n$ can be known by a combinatorial or other argument and furthermore the graph has a high symmetry with its local parameters easily determined by inspection, one may estimate the diameter using an algebraic, geometrical, or probabilistic argument. In our previous study,Hirata (2024) we proposed a probabilistic estimate of the diameter for a large Cayley graph, which reproduces the diameters of the Rubik’s Cube groups within 2 of the actual values or often exactly. Here, we introduced a strict lower bound for the diameter for a symmetric Cayley graph, calculable with $n$ and other easily determined local parameters such as the degree $k$, girth $g$, and number $\eta$ of $g$-cycles per vertex. Similar, but more crude estimates of the diameters of the Rubik’s Cube groups have been made,Singmaster (1981) but this study is distinguished from them for presenting a rigorous lower bound also applicable to other symmetric graphs. When applied to symmetric Cayley graphs of some Rubik’s Cube groups, it yields tighter lower bounds for the diameters than those of random $k$-regular graphs proposed by Bollobás and de la Vega,Bollobás and de la Vega (1982); Bollobás (2001) which are 60% to 77% of the correct values for large-$n$ graphs.